OpenStudy (anonymous):

Add. Express your answer in simplest terms. 2/3 + 2/7

OpenStudy (anonymous):

You have any idea how to do it?

OpenStudy (anonymous):

find the LCD

OpenStudy (mathmale):

Yes. What is the LCD?

OpenStudy (anonymous):

Well it doesn't have to be the lowest common denominator, just a common denominator, although the LCD is the most convenient.

OpenStudy (anonymous):

Wanna know the easiest way of finding a common denominator?

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

K. \[\large \sf For~\frac{a}{b}+\frac{c}{d}\\\large \sf (\frac{a}{b} \times \frac{d}{d})+(\frac{c}{d} \times \frac{b}{b})\] Keep in mind that this will just get you a common denominator and it may not be the lowest one, but it is a good method to keep in hand

OpenStudy (mathmale):

Just to be clear: the LCD here is the product bd. In the actual addtiion problem you're working on, \[\frac{ 2 }{ 3 }+\frac{ 3 }{ 7, }\]

OpenStudy (anonymous):

2/7*

OpenStudy (mathmale):

the two diferent denominators are 3 and 7. Multiply those together to get y our LCD. What is it? Yes, 2/7, thank you.

OpenStudy (mathmale):

\[\frac{ 2 }{ 3 }+\frac{ 2 }{ 7,}\]

OpenStudy (mathmale):

@dulcereon, please multiply the two denominators together to obtain the LCD.

OpenStudy (anonymous):

\[\large \sf For~\frac{a}{b}+\frac{c}{d}\] \[\large \sf \huge ( \large \frac{a}{b} \times \frac{d}{d} \huge ) \large + \huge ( \large \frac{c}{d} \times \frac{b}{b} \huge )\]

OpenStudy (anonymous):

Made the parentheses less ugly

OpenStudy (mathmale):

@dulceceron, please multiply the two denominators together to obtain the LCD. Where have you been?

OpenStudy (anonymous):

the LCD is 21 and do i x or add the numerators

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

\[\large \sf For~\frac{a}{b}+\frac{c}{d}\] \[\large \sf \huge ( \large \frac{a}{b} \times \frac{d}{d} \huge ) \large + \huge ( \large \frac{c}{d} \times \frac{b}{b} \huge )\] for a=2, b=3, c=2, d=7

OpenStudy (anonymous):

i dont get the letters

OpenStudy (mathmale):

You definitely MULTIPLY the 2 denominators together. By what number do you multiply the 3 in the first denominator?

OpenStudy (anonymous):

They're just representations of the fractions so you can use the same process on other fractions.

OpenStudy (mathmale):

The Sadist is just taking a more theoreticl approach, but is entirely correct.

OpenStudy (mathmale):

You definitely MULTIPLY the 2 denominators together. By what number do you multiply the 3 in the first denominator?

OpenStudy (anonymous):

do i multiply the numerators

OpenStudy (mathmale):

Yes. Both The Sadist and I have asked you to do that.

OpenStudy (mathmale):

The left fraction is 2/3. I'm asking you: By what number do you multiply that 3, the denominator of the left fraction?

OpenStudy (anonymous):

4/21

OpenStudy (mathmale):

The 21 is fine; that's your LCD. Where did that 4 come from? We were not yet ready to combine the two fractions. The left fraction is 2/3. By which number do we multiply that 3 to get 27?

OpenStudy (anonymous):

si

OpenStudy (mathmale):

si? as in Spanish, si? Please answer with a number.

OpenStudy (anonymous):

yes and what do you mean answer with a number

OpenStudy (mathmale):

You have the fraction 2/3. You must find the LCD to add 2/3 to 2/7. You correctly found that the LCD is 21. How did you find that 21?

OpenStudy (mathmale):

3 ( ? ) = 21

OpenStudy (anonymous):

can u plz tell me the answer cuz i am in a hurry

OpenStudy (mathmale):

3 (7) = 21. the first fraction is 2/3. I'm asking you to multiply BOTH the numerator (2) and the denominator (3) by 7. 7(2) = 7(3) =

OpenStudy (anonymous):

6/7

OpenStudy (mathmale):

but 7*2 = 14. where did 6/7 come from? 7*3= what?

OpenStudy (anonymous):

21

OpenStudy (mathmale):

If this multiplication is hard for you, use a calculator. But you'll need to know how to multiply 2 by 7 and 3 by 7.

OpenStudy (anonymous):

just tell me the answer

OpenStudy (mathmale):

Sorry, I cannot and will not do that. My purpose, and that of OpenStudy, is to help YOU learn how to do these problems YOURSELF.

OpenStudy (mathmale):

2 times 7 is 2*7 or 2(7) and is 14. 3 times 7 is 3*7, and is .... ??

OpenStudy (anonymous):

i found the answer

OpenStudy (mathmale):

Good for you, but I wish so much you'd take the time to learn how to do this multiplication y ourself.

OpenStudy (mathmale):

Good luck. Come back when you're feeling more ready to learn and practice.